# Floquet Engineering of Relativistic Electrons using Propagating Waves

### v1

@@ Abstract
This study examines the electronic states of quantum materials influenced by waves $V(Qx-\omega t)$ propagating through them. These waves can include laser light, polaritons, acoustic waves, and sliding density waves, whose speed may vary widely. Since these waves are periodic in both space and time, the electrons in the materials form Floquet-Bloch states when influenced by the waves. We investigate band engineering by these waves, using three-dimensional (3D) Dirac electrons in circularly polarized laser wave as a case study. We classify the fields by their speed into time, light and space-like waves, and find that the electronic states are distinctive in the three cases. We demonstrate phase transitions of the electronic bands and discover emergent Floquet Weyl bands, which potentially become Type-II in the vicinity of the transition. The effect of Floquet band engineering is enhanced by Lorentz contraction, which is maximized when the wave speed approaches the Fermi velocity. Additionally, we discuss the geometric nature of the response currents.
@@
Introduction.— Floquet engineering, i.e., a method to coherently control system using carefully designed time-periodic driving fields, has been applied to various systems to realize exotic states of matter Eckardt (2017); Oka and Kitamura (2019); Rudner and Lindner (2020); de la Torre et al. (2021); Ozawa et al. (2019). This technique enables the study of Floquet topological insulators Oka and Aoki (2009); Lindner et al. (2011); Kitagawa et al. (2011), anomalous Floquet topological insulators Rudner et al. (2013), and Floquet Weyl semimetals Wang et al. (2014); Ebihara et al. (2016); Chan et al. (2016); Hübener et al. (2017); Yoshikawa et al. (2022); Bucciantini et al. (2017); Hirai et al. (2024) as well as other exotic many-body states Harper et al. (2020); Mori (2023). While many applications assume spatially homogeneous driving fields, Floquet theory can also be applied to spatially varying drives. A propagating wave $V(Qx-\omega t)$ moving in the $x$-direction with speed $v=\omega/Q$ exemplifies this. There are numerous instances of propagating waves in quantum materials where their interaction with electrons presents intriguing challenges.These include laser light (mostly treated as homogeneous fields Wang et al. (2013); McIver et al. (2020); Ito et al. (2023)), slow light using photonic crystals Baba (2008), sliding density waves Grüner (1988); Lee et al. (1973); Monçeau et al. (1976); Bardeen (1979); Cox et al. (2008), surface acoustic waves Puebla et al. (2020), and polaritons Basov et al. (2021); Walicki et al. (2024a). The coherence and high intensity of these waves make the nonperturbative response, explored through Floquet theory Yang et al. (2024); Walicki et al. (2024b), particularly interesting.
The nonlinear response of quantum materials to external fields, especially considering geometric effects, has become a vibrant research area. Studies employing Floquet theory M. Torres and A. Kunold (2005); Oka and Aoki (2009, 2011), perturbation theory Sodemann and Fu (2015); Sipe and Shkrebtii (2000); Morimoto and Nagaosa (2016); Ahn and Nagaosa (2021); Ahn et al. (2022), and tunneling theory Takayoshi et al. (2021a); Kitamura et al. (2020) have been conducted. Generalizing the linear response theory of electric conductivity to the non-perturbative domain through propagating waves offers another pathway. In linear response theory Kubo (1957), we consider the current ($i,j=x,y,z$)
$J_{i}(\bm{r},t)=\sum_{j}\sigma_{ij}(\bm{Q},\omega)E_{j}(\bm{r},t),$ (1)
induced by an applied ac-electric field $E_{i}(\bm{r},t)=E_{i}(\bm{Q},\omega)e^{i(\bm{Q}\cdot\bm{r}-\omega t)}$. The zero-temperature conductivity $\sigma_{ij}(\bm{Q},\omega)$ is obtained by evolving the ground state |$\ket{\psi}$ under the field $V(t)$ turned on in the infinite past:
[…]
Here, $V(t)$ represents the ac-electric field. The Kubo formula, obtained by evaluating $\bra{\psi^{\prime}}j_{i}\ket{\psi^{\prime}}$ assuming adiabatic connection of |$\ket{\psi^{\prime}}$ to |$\ket{\psi}$, is expressed as a perturbative series in $V$. Several questions arise: (a) When is the perturbative picture of |$\ket{\psi^{\prime}}$ applicable? (b) In the calculation of the dc-conductivity, we take the limit $\bm{Q}\to\bm{0}$ and $\omega\to 0$ in this order. What happens if we consider the reverse order? (c) What are the new geometric properties that emerge in the nonperturbative regime? Applying Floquet theory, we can non-perturbatively study the properties of the state in an propagating wave Eqn. (2) and answer these questions. To proceed with a concrete example, this work focuses on three-dimensional (3D) Dirac electrons in propagating waves. Dirac electrons are relativisitic and its Lorentz covariance enables us to categorize the propagating waves into time, space and light-like ones according to its speed relative to the Fermi velocity. The resulting Floquet-Bloch states have distinctive properties in each of these cases, leading to unique electronic behaviors and potential phase transitions.
@@
Floquet-Bloch theorem for electrons in propagating waves.— Let us study non-interacting three dimensional (3D) Dirac electrons  Yan and Felser (2017); Gorbar et al. (2021); Fedotov et al. (2023) in a propagating wave as a prototypical example. We employ the four vector notation of relativity theory, which is useful in non-relativistic system as well. The space-time and momentum-energy variables are expressed as $x^{\mu}=(ct,\bm{x})$ and $k^{\mu}=(\varepsilon/c,\bm{k})$, and their product is $k\cdot x=k^{\mu}x_{\mu}=\varepsilon t-\bm{k}\cdot\bm{x}$. We use Einstein’s summation convention and $v_{\mu}=\eta_{\mu\nu}v^{\nu}$ with metric $\eta_{\mu\nu}=\mbox{diag}(1,-1,-1,-1)$. The speed of light $c$ is introduced as the Fermi velocity of the Dirac electron. We consider a Lagrangian (see e.g., Ref. Bjorken and Drell (1964))
@@
where the Floquet-Bloch state has a periodicity of $q\cdot x\sim q\cdot x+2\pi$ ($N$ is a normalization factor). Note that this periodicity includes the spatial $x\sim x+2\pi/Q$ and temporal $T\sim T+2\pi/\omega$ periodicities when $Q\neq 0$ and $\omega\neq 0$ respectively. For the Floquet-Bloch state (Eqn. (7)) to be a solution of the time dependent Schrödinger equation, the Floquet quasi-momentum-energy vector $k^{\mu}=(\varepsilon/c,\bm{k})$ and the state |$\ket{\Phi}$ must satisfy the Floquet-Bloch equation. We employ Sambe’s space-time picture denoting the Floquet-Bloch state as an infinite dimensional vector with Fourier modes as components |$\ket{\Phi_{\bm{k}}}=^{t}(\ldots,\ket{\phi^{n-1}_{\bm{k}}},\ket{\phi^{n}_{\bm{k}}},\ket{\phi^{n+1}_{\bm{k}}},\ldots)$. The Schrödinger equation can be expressed as an eigenvalue problem that determines the $\alpha$-th eigenstate Yang et al. (2024),
@@
This equation can be considered as an extension of the standard Floquet equation, with the $ncq_{0}$-term, and Bloch equation, with the $+n\bm{q}$ term, mixing them into a single defining equation of the Floquet-Bloch state.
The geometric properties of this state can be captured by the Berry connection four vector $\mathcal{A}_{\alpha\bm{k}}^{\mu}=(c\mathcal{A}_{\alpha\bm{k}}^{0},\mathcal{A}_{\alpha\bm{k}}^{i})$ of the $\alpha$-th state, whose components are defined by
@@
This result shows that the longitudinal current $J^{x}_{\rm tot}(Qx-\omega t)=\frac{d}{dt}P^{x}_{\rm tot}(Qx-\omega t)$ that flows parallel to the propagating wave $V(x)$ is non-dissipative and is a geometric quantity. This current can be considered as one of the non-perturbative generalization of the Kubo formula Eqn. (1), in particular, the quantum Hall current Thouless et al. (1982); Kohmoto (1985). This can be noticed if we consider an electric field in the $y$-direction $E_{y}(Q,\omega)$ as our driving field by setting $V=\frac{E_{y}(Q,\omega)}{\omega}\Gamma^{y}$. Then, the current $J^{x}_{\rm tot}(Qx-\omega t)$ is a non-perturbative Hall current that flows in the direction perpendicular to the electric field.
Application to relativistic fermions.— Below, we focus our discussion to applications on Dirac electrons. Although we can directly study the Floquet-Bloch states in waves with general $q^{\mu}$-vector by solving Eqn. (8), we can also use the Lorentz covariance of the Dirac electron to study the band engineering systematically. To this end, we perform a coordinate transformation to a “rest frame” in which the wave is stopped.
@@
We set the “speed of light” parameter used in the above formulation to be the Fermi velocity of the Dirac electron, i.e., $c=v_{F}$. In contrast to elementary particle physics, in condensed matter systems we can consider waves in materials faster than the Dirac electrons. Let us call waves that propagate faster than the Fermi velocity as time-like ($v/v_{F}>1$), equally fast as light-like ($v/v_{F}=1$), and those that are slower as space-like ($v/v_{F}<1$). We expect enhancement of relativistic effects when the speed of the propagating wave $v$ approaches $v_{F}$. We introduce a Lorentz boost $x^{\prime}=ax$ (or $x^{\prime\mu}=a^{\mu}{}_{\nu}x^{\nu}$) given by
@@
Next, we relate the electron wave function in the two frames using Lorentz transformation. To this end, we enforce the fields to be related by Lorentz transformation
@@
with $A$ determining its strength. Physically, this wave represents circularly polarized laser fields propagating in the $x$-direction with $A=E/\omega$ related to the electric field $E$. This wave is invariant under Lorentz boost in the $x$-direction and thus $V^{\prime}(x^{\prime})=V(ax)$. Below, we assume that we can select its speed $v$ freely.
@@
Massive Dirac electron and phase transitions.— Next, let us study the case of a massive Dirac electron in the propagating wave Eqn. (28). We focus on the low energy Floquet-Bloch bands that descend from the upper and lower Dirac bands. In the time-like case, we expect the spectrum at large $\omega$ to be given by
@@
in the rest frame Ebihara et al. (2016). The gap closes when the strength of the chiral gauge field exceeds the mass parameter $A_{5{\rm eff}}^{\prime x}=m$. In Fig. 3 (i)-(iv), we depict the Floquet-Bloch spectrum for several values of $\omega$ while fixing the wave number $Q=10$. In the case of large $\omega$ as in Fig. 3 (i), the upper and lower Dirac bands split by the chiral gauge field and the gap at $\bm{k}=0$ becomes small. A phase transition occur to a state with two Weyl fermions as depicted in Fig. 3 (ii) when the gap close at $A_{5{\rm eff}}^{\prime x}=m$. Near the phase transition, the Weyl fermions are type-II Soluyanov et al. (2015) since the bands are tilted by the Lorentz boost. When we further approach the light-like limit $\omega\sim Q$, the energy spectrum becomes dense since the Brillouin zone becomes narrow and multiple band foldings occur. This is the region where the system approaches the adiabatic limit $\omega^{\prime}\to 0$ in the rest frame. In this limit, we expect adiabatic time evolution with nonperturbative geometric effects dominates the dynamics Takayoshi et al. (2021b). In the space like case, the gap is always open, while the bands are tilted and shows a chiral flow (Fig. 3 (iv)). Weyl points can exist if $A_{5{\rm eff}}^{\prime 0}>m$ is satisfied in the rest frame.
Discussion.— We have studied the Floquet-Bloch band engineering of Dirac fermions induced by propagating waves. The Lorentz covariance of the Dirac equation allows us to classify the waves into three categories: time-like, light-like, and space-like, and the corresponding electronic bands are described by Floquet states, Volkov states Volkov (1935), and Bloch states, respectively. One of the most striking findings is that the Floquet-Bloch bands undergo significant modifications near the light-like case due to the Lorentz contraction factor $\gamma$ appearing in the effective Floquet terms as in Eqns. (29, 30). The ability to design propagating waves with specific characteristics could enable efficient realization of Floquet engineering of band topology in quantum materials.
@@ ACKNOWLEDGMENTS
The authors appreciate the fruitful discussions with S. Kitamura, S. Chaudhary, and K. Saha. This work is supported by JSPS KAKENHI (No. JP23H04865, No. JP23K25837 and No. JP23K22487), MEXT, Japan, and JST CREST Grant No. JPMJCR19T3, Japan.

### v2

# Relativistic Floquet States Near the Shockwave Limit

@@ Abstract
We investigate the impact of propagating waves $V(Qx-\omega t)$ on the electronic states of quantum materials, focusing on Floquet-Bloch states of relativistic fermions. These waves include laser light, polaritons, acoustic waves, and sliding density waves, each propagating at different speeds. Generically, we categorize the fields based on their speed relative to the Fermi velocity into time-like, light-like, and space-like waves, finding that the resulting electronic states exhibit distinct characteristics in each cases. For 3D Dirac electrons in circularly polarized waves, we identify emergent Floquet Weyl bands, which can transition into Type-II Weyl states near a critical point. The effects of Floquet band engineering are significantly enhanced by Lorentz contraction, which reaches its maximum when the wave speed approaches but slightly exceeds the Fermi velocity, leading to the realization of a shockwave-like state. The shockwave-like states are related to the near adiabatic dynamics via Lorentz transformation.
@@
Introduction.— Floquet engineering enables the coherent control of quantum systems through carefully designed time-periodic driving fieldsEckardt (2017); Oka and Kitamura (2019); Rudner and Lindner (2020); de la Torre et al. (2021); Ozawa et al. (2019). This approach has been widely applied to study Floquet topological insulators Oka and Aoki (2009); Lindner et al. (2011); Kitagawa et al. (2011), anomalous Floquet topological insulators Rudner et al. (2013), and Floquet Weyl semimetals Wang et al. (2014); Ebihara et al. (2016); Chan et al. (2016); Hübener et al. (2017); Yoshikawa et al. (2022); Bucciantini et al. (2017); Hirai et al. (2024), as well as other exotic phenomena Harper et al. (2020); Mori (2023). While most applications assume spatially homogeneous driving fields, Floquet theory can also be extended to spatially varying drives. A propagating wave $V(Qx-\omega t)$ moving along the $x$-direction serves as a prime example. Many types of propagating waves exist in quantum materials, and their interaction with electrons can be harnessed for Floquet engineering of material properties. The primary tool for Floquet engineering has been laser light Wang et al. (2013); McIver et al. (2020); Ito et al. (2023); Yoshikawa et al. (2022), while its speed can be significantly reduced in specific experimental setups Baba (2008). Other examples include sliding density waves Grüner (1988); Lee et al. (1973); Monçeau et al. (1976); Bardeen (1979); Cox et al. (2008), surface acoustic waves Oliva-Leyva and Naumis (2015); Puebla et al. (2020), and polaritons Basov et al. (2021); Walicki et al. (2024a). The coherence and high intensity of these waves make the nonperturbative response, explored through Floquet theory Yang et al. (2024); Walicki et al. (2024b), particularly interesting.
The nonlinear response of quantum materials hosting relativistic quasiparticles, such as Dirac and Weyl electrons, has been an area of active research Wang et al. (2014); Ebihara et al. (2016); Chan et al. (2016); Hübener et al. (2017); Yoshikawa et al. (2022); Bucciantini et al. (2017); Hirai et al. (2024); Ahn and Nagaosa (2021); Ahn et al. (2022); Takayoshi et al. (2021a). Various theoretical approaches, including Floquet theory M. Torres and A. Kunold (2005); Oka and Aoki (2009, 2011), perturbation theory Sodemann and Fu (2015); Sipe and Shkrebtii (2000); Morimoto and Nagaosa (2016); Ahn and Nagaosa (2021); Ahn et al. (2022), and tunneling theory Takayoshi et al. (2021a), have been employed to investigate these systems. Let $v_{F}$ denote the speed of the relativistic quasiparticles (i.e., the Fermi velocity). Consider a scenario where these particles exist in the presence of a coherent propagating wave $V(Qx-\omega t)$ traveling at speed $v=\omega/Q$ What happens when $v$ approaches and slightly exceeds $v_{F}$? Classical wave theory predicts that a shockwave of quasiparticles should form. Here, we explore the quantum mechanical counterpart of this phenomenon and investigate its fundamental nature. To address this question, we extend Floquet theory to a Floquet-Bloch framework for propagating waves Yang et al. (2024). As a concrete example, we focus on three-dimensional (3D) Dirac electrons subjected to propagating waves. The Lorentz covariance of the problem allows us to classify the waves into time-like, space-like, and light-like categories based on their speed relative to the Fermi velocity. The resulting Floquet-Bloch states exhibit distinct properties in each case, leading to unique electronic behaviors and potential phase transitions.
Figure 1: (a) In the Floquet-Bloch approach, the bands folds in the space $(\varepsilon,k)$ of Floquet pseudo-energy $\varepsilon$ and crystal-momentum $k$, and the folding direction is given by a space-time reciprocal vector $q^{\mu}=(\omega/c,\bm{Q})$. (b) Floquet-Bloch spectrum for a 1D Dirac electron obtained as the eigenvalues of the Floquet-Bloch Hamiltonian Eq. (6) with $V_{0}=0.2,\;Q=1.5,\;\omega=0.7$.
@@
Floquet-Bloch theorem for electrons in propagating waves.— Let us study non-interacting three dimensional (3D) Dirac electrons  Yan and Felser (2017); Gorbar et al. (2021); Fedotov et al. (2023) in a propagating wave as a prototypical example. We employ the four vector notation of relativity theory. However, we note that the formalism is valid for non-relativistic system as well. The space-time and momentum-energy variables are expressed as $x^{\mu}=(ct,\bm{x})$ and $k^{\mu}=(\varepsilon/c,\bm{k})$, and their product is $k\cdot x=k^{\mu}x_{\mu}=\varepsilon t-\bm{k}\cdot\bm{x}$. We use Einstein’s summation convention and $v_{\mu}=\eta_{\mu\nu}v^{\nu}$ with metric $\eta_{\mu\nu}=\mbox{diag}(1,-1,-1,-1)$. The speed of light $c$ is introduced as a book keeping factor. We consider a Lagrangian (see e.g., Ref. Bjorken and Drell (1964))
@@
where the Floquet-Bloch state has a periodicity of $q\cdot x\sim q\cdot x+2\pi$ ($N$ is a normalization factor). Note that this periodicity includes the spatial $x\sim x+2\pi/Q$ and temporal $T\sim T+2\pi/\omega$ periodicities when $Q\neq 0$ and $\omega\neq 0$ respectively. For the Floquet-Bloch state (Eqn. (5)) to be a solution of the time dependent Schrödinger equation, the Floquet quasi-momentum-energy vector $k^{\mu}=(\varepsilon/c,\bm{k})$ and the state |$\ket{\Phi}$ must satisfy the Floquet-Bloch equation. We employ Sambe’s space-time picture denoting the Floquet-Bloch state as an infinite dimensional vector with Fourier modes as components |$\ket{\Phi_{\bm{k}}}=^{t}(\ldots,\ket{\phi^{n-1}_{\bm{k}}},\ket{\phi^{n}_{\bm{k}}},\ket{\phi^{n+1}_{\bm{k}}},\ldots)$. The Schrödinger equation can be expressed as an eigenvalue problem that determines the $\alpha$-th eigenstate,
@@
This equation can be considered as an extension of the Floquet equation (with a $ncq_{0}$-term) and the Bloch equation(with a $n\bm{q}$-term), mixing them into a single defining equation for the Floquet-Bloch state. In Fig. 1 (a), we show how the replica side bands are formed from the original band by shifting it with the reciprocal vector $q^{\mu}$. Figure 1 (b) plots the spectrum for a 1D Dirac electron $H(k_{x})=\sigma_{x}k_{x}$ in a wave $V(x)=V_{0}\sigma_{y}\cos(Qx-\omega t)$.
Geometric aspect of the Floquet-Bloch states in propagating waves.— The formulation of the geometric properties of the Floquet-Bloch states follows those of the Bloch states Xiao et al. (2010). We can define the Berry connection four vector $\mathcal{A}_{\alpha\bm{k}}^{\mu}=(c\mathcal{A}_{\alpha\bm{k}}^{0},\mathcal{A}_{\alpha\bm{k}}^{i})$ of the $\alpha$-th state, whose components are defined by
@@
This result shows that the longitudinal polarization current $J^{x}_{\rm tot}(Qx-\omega t)=\frac{d}{dt}P^{x}_{\rm tot}(Qx-\omega t)$ that flows parallel to the propagating wave $V(x)$ is non-dissipative and is a geometric quantity. This current can be considered as one of the non-perturbative generalization of the Kubo formula, in particular, the quantum Hall current Thouless et al. (1982); Kohmoto (1985). This can be noticed if we consider an electric field in the $y$-direction $E_{y}(Q,\omega)$ as our driving field by setting $V=\frac{E_{y}(Q,\omega)}{\omega}\Gamma^{y}$. Then, the current $J^{x}_{\rm tot}(Qx-\omega t)$ is a non-perturbative Hall current that flows in the direction perpendicular to the electric field.
Application to relativistic fermions.— Below, we focus our discussion to applications on Dirac electrons. Although we can directly study the Floquet-Bloch states in waves with general $q^{\mu}$-vector by solving Eqn. (6), we can also use the Lorentz covariance of the Dirac electron to study the states systematically. To this end, we perform a coordinate transformation to a “rest frame” in which the wave becomes stopped or homogeneous.
@@
We set the “speed of light” parameter used in the above formulation to be the Fermi velocity of the Dirac electron, i.e., $c=v_{F}$. Let us call waves that propagate faster than the Fermi velocity as time-like ($v/v_{F}>1$), equally fast as light-like ($v/v_{F}=1$), and those that are slower as space-like ($v/v_{F}<1$). We expect enhancement of relativistic effects when the speed of the propagating wave $v$ approaches $v_{F}$. We introduce a Lorentz boost $x^{\prime}=ax$ (or $x^{\prime\mu}=a^{\mu}{}_{\nu}x^{\nu}$) given by
@@
Next, we relate the wave function of the electrons in the two frames using Lorentz transformation. To this end, we enforce the fields to be related by Lorentz transformation
@@
with $A$ determining its strength. Physically, this wave represents circularly polarized laser fields propagating in the $x$-direction with speed $v=\omega/Q$ with a field parameter $A=E/\omega$ with $E$ being the electric field strength. This wave is invariant under Lorentz boost in the $x$-direction and thus $V^{\prime}(x^{\prime})=V(ax)$. Below, we assume that we can select its speed $v$ freely.
@@
Massive Dirac electron and phase transitions.— Finally, let us study the case of a massive Dirac electron in the propagating wave Eqn. (26). We focus on the low energy Floquet-Bloch bands that descend from the upper and lower Dirac bands. In the time-like case, we expect the spectrum at large $\omega$ to be given by
@@
in the rest frame Ebihara et al. (2016). The gap closes when the strength of the chiral gauge field exceeds the mass parameter $A_{5{\rm eff}}^{\prime x}=m$. In Fig. 4 (i)-(iv), we depict the Floquet-Bloch spectrum for several values of $\omega$ while fixing the wave number $Q=10$. In the case of large $\omega$ as in Fig. 4 (i), the upper and lower Dirac bands split by the chiral gauge field and the gap at $\bm{k}=0$ becomes small. A phase transition occur to a state with two Weyl fermions as depicted in Fig. 4 (ii) as the gap close at $A_{5{\rm eff}}^{\prime x}=m$. Near the phase transition, the Weyl fermions are type-II Soluyanov et al. (2015) since the bands are tilted by the Lorentz boost. When we further approach the light-like limit $\omega\sim Q$, the energy spectrum becomes dense since the Brillouin zone becomes narrow and multiple band foldings occur. This is the region where the system approaches the shockwave limit $v\to v_{F}$. In fact, the states realized in the shockwave limit are related to the near adiabatic states at $\omega^{\prime}\to 0$ in the rest frame by Lorentz transformation. In this limit, we expect adiabatic time evolution with nonperturbative geometric effects dominates the dynamics Takayoshi et al. (2021b). On the other hand, in the space like case, the gap is always open, while the bands are tilted and shows a chiral flow (Fig. 4 (iv)). Weyl points can exist if $A_{5{\rm eff}}^{\prime 0}>m$ is satisfied in the rest frame.
Discussion.— We have studied the Floquet-Bloch band engineering of Dirac fermions induced by propagating waves. The Lorentz covariance of the Dirac equation allows us to classify the waves into three categories: time-like, light-like, and space-like, and the corresponding electronic bands are described by Floquet states, Volkov states Volkov (1935), and Bloch states, respectively. One of the most striking findings is that the Floquet-Bloch bands undergo significant modifications near the light-like case due to the Lorentz contraction factor $\gamma$ appearing in the effective Floquet terms as in Eqns. (27, 28). The ability to design propagating waves with specific characteristics could enable efficient realization of Floquet engineering of band topology in quantum materials.
@@ ACKNOWLEDGMENTS
The authors appreciate the fruitful discussions with Christopher Yang, Iliya Esin, Sota Kitamura, Swati Chaudhary, and Kush Saha. This work is supported by JSPS KAKENHI (No. JP23H04865, No. JP23K25837 and No. JP23K22487), MEXT, Japan, and JST CREST Grant No. JPMJCR19T3, Japan.